OpenStudy (mbma526):

Select the equation that contains the point (4, -8), and in which the graph of the line has a positive slope y – 8 = –4(x + 4) y – 8 = 4(x + 4)** i thought it was this one y + 8 = +4(x – 4) y + 8 = –4(x – 4)

OpenStudy (welshfella):

a positive slope has a positive coefficient of x.

OpenStudy (welshfella):

so its i either B or C why do you think its B?

OpenStudy (mbma526):

It seems like the one that would make since

OpenStudy (welshfella):

the general form is y - y1 = m(x - x1)

OpenStudy (welshfella):

Dont forget (x1,y1) is the point on the line. So , here, x1 = 4 and y1 = -8 Plug these into the general equation and see what you get.

OpenStudy (mbma526):

I got C

OpenStudy (mbma526):

@welshfella

OpenStudy (welshfella):

yes C is correct

OpenStudy (mbma526):

Thank you @welshfella

OpenStudy (welshfella):

y - (-8) - 4(x - 4) y +8 = 4(x - 4) yw